OpenStudy (anonymous):

MEDAL TO WHOEVER CAN ANSWER! A Web music store offers two versions of a popular song. The size of the standard version is megabytes (MB). The size of the high-quality version is MB. Yesterday, there were downloads of the song, for a total download size of MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

Do you have any numbers that is given?

OpenStudy (anonymous):

lol yeah sorry that would be helpful huh?

OpenStudy (anonymous):

Yes ma'am, it would.

OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.1 megabytes (MB). The size of the high-quality version is 4.2 MB. Yesterday, there were 1300 downloads of the song, for a total download size of 4578 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

There ya go! Sorry for that lol

OpenStudy (anonymous):

Okay, start solving this and tell me what you got. 2.2(1400−y)+4.7y=5105

OpenStudy (anonymous):

2.1 I mean

OpenStudy (anonymous):

Wait, are you solving at the moment?

OpenStudy (anonymous):

I couldn't figure out how to set it up, it's been a while since I've done word problems. As long as you could give me the equation I could solve. lol

OpenStudy (anonymous):

but yes, I'm doing so lol

OpenStudy (anonymous):

I messed up the equation. Solve this: 2.1(1300-y)+4.2y=4578

OpenStudy (anonymous):

haha, no worries and I figured that's what you meant :P

OpenStudy (anonymous):

@sourwing stay here for a while - in case I don't get it right.

OpenStudy (anonymous):

is it y=880?

OpenStudy (anonymous):

Give me a minute to solve it as well.

OpenStudy (anonymous):

Okay (:

OpenStudy (anonymous):

Yes, you're correct!

OpenStudy (anonymous):

See, I'm not totally stupid, lol I just have trouble coming up with the equation, I can solve no problem :P and thank you soooo much! :D

OpenStudy (anonymous):

Now solve this equation x+880=1300

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

420

OpenStudy (anonymous):

So far we know x = 420 y = 880

OpenStudy (anonymous):

So now what do we do?

OpenStudy (anonymous):

You found both answers. Correct?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

That's how many downloads of the standard version.

OpenStudy (anonymous):

420 is the answer?

OpenStudy (anonymous):

cool :D

OpenStudy (anonymous):

Yes, 420 for x, and y is 880. But make sure to show these two answers for your teacher.

OpenStudy (anonymous):

@sourwing I guess I didn't need you after all. Sorry to keep you waiting lol.

OpenStudy (anonymous):

Okay, thank you very much!

OpenStudy (anonymous):

Anytime, you did great yourself :)